# Lyapunov Dynamics in Entangled Biphoton Spectroscopy

### v1

@@ IV Summary
Recent advancements in quantum light spectroscopy have revealed exciting possibilities of using entangled photons as sensitive probes of many-body dynamics and correlations in materials. Although experimental and theoretical efforts continue to evolve in parallel, the literature still lacks a framework that accurately explains observed phenomena. The exponential growth of the Hilbert space and the presence of nonlinear and many-body effects make direct simulations computationally prohibitive. To address this, we reformulated the problem using a finite-sized correlation matrix based on a bilinear bosonic Hamiltonian leading to gaussian-preserving time evolutions. This approach allows us to retain the essential physical dynamics that occurring in bipolaritonic systems.
We applied our model to an experimentally measured joint spectral intensity (JSI) of frequency-entangled biphotons generated through spontaneous parametric down conversion. Our simulations predict a shift in the output JSI that is consistent with the experimental observations. Moreover, we observed the emergence of off-diagonal peaks in the JSI, which indicate enhanced correlations between specific signal-idler frequency pairs. We interpret these features as signatures of a discrete cavity mode decaying into the continua of signal and idler modes. This decay redistributes the spectral amplitudes and gives rise to interference effects similar to Fano resonances.
@@
V Acknowledgments
The work at the University of Houston was funded by the National Science Foundation (CHE-2404788) and the Robert A. Welch Foundation (E-1337). ARSK acknowledges funding from the National Science Foundation CAREER grant (CHE-2338663), start-up funds from Wake Forest University, funding from the Center for Functional Materials at Wake Forest University.
VI Data Availability Statement

### v2

@@ IV Summary
Recent advancements in quantum light spectroscopy have revealed exciting possibilities of using entangled photons as sensitive probes of many-body dynamics and correlations in materials. Although experimental and theoretical efforts continue to evolve in parallel, the literature still lacks a framework that accurately explains observed phenomena. The exponential growth of the Hilbert space and the presence of nonlinear and many-body effects make direct simulations computationally prohibitive. To address this, we reformulated the problem using a finite-sized correlation matrix based on a bilinear bosonic Hamiltonian, leading to Gaussian-preserving time evolutions. This approach allows us to retain the essential physical dynamics that occur in bipolaritonic systems.
We applied our model to an experimentally measured JSI of frequency-entangled biphotons generated through spontaneous parametric down conversion. Our simulations predict a shift in the output JSI that is consistent with the experimental observations. Moreover, we observed the emergence of off-diagonal peaks in the JSI, which indicate enhanced correlations between specific signal-idler frequency pairs. We interpret these features as signatures of a discrete cavity mode decaying into the continua of signal and idler modes. This decay redistributes the spectral amplitudes and gives rise to interference effects similar to Fano resonances.
@@
Acknowledgements.
The work at the University of Houston was funded by the National Science Foundation (CHE-2404788) and the Robert A. Welch Foundation (E-1337). ARSK acknowledges funding from the National Science Foundation CAREER grant (CHE-2338663), start-up funds from Wake Forest University, funding from the Center for Functional Materials at Wake Forest University.Author Contributions
S.D. developed the theoretical framework, performed the simulations, and wrote the manuscript. A.R.S.K. and E.R.B. supervised the project and contributed to discussions and manuscript revisions. All authors reviewed and approved the final version of the manuscript.
Data Availability Statement